A bicyclist races once around a circular track of radius 265 m every 2.9 minutes at a constant speed. What is the bicyclist's acceleration in units of meters/second2?
Answer #1

Solna 278 distance) Linear speed CV mils: 2.94 Go (time) V = 243-144 265 mis. 2.9x60 V = 9.56 44 m/s Acceleration = ( centar